The Addition Property of Equality:
Solve the problem.
5. Is k = -2 a solution of 5k – 6(k – 1) = 8?

Answer by algebrahouse.com(1659)

Is k = -2 a solution of 5k – 6(k – 1) = 8?

Substitute -2, in for k, into the equation.
If you get a true statement, then it is a solution.
If you get a false statement, then it is not a solution.

5(-2) - 6(-2 - 1) = 8 {substituted -2 in for k}
-10 - 6(-3) = 8 {multiplied 5 by -2 and subtracted -2 - 1}
-10 + 18 = 8 {multiplied -6 by -3}
8 = 8 {added -10 and 18}
- this is a true statement

-2 is a solution of the equation

For more help from me, visit: www.algebrahouse.com